1. Glenn collected 3 more bags of bottles for recycling than Paula did. They collected a total of 25 bags of bottles for recycling. How many bags did Paula collect? The domain of the solution is {0, 4, 11, 22}.
Answer by macston(5194)

.
P=bags Paula collected; G=bags Glenn collected=P+3 bags
G+P=25 bags Substitute for G
(P+3 bags)+P=25 bags Subtract 3 bags from each side.
2P=22 bags Divide each side by 2.
P=11 bags ANSWER: Paula collected 11 bags.